# Risk-Calibrated Human-Robot Interaction via Set-Valued Intent Prediction

### v1

@@ III-C Goal: Certifiable Autonomy
Situational ambiguity results in many potentially correct robot actions arising with potentially no safe external resolution, save for direct human intervention (see e.g. Fig. 1). Our goal in this work is to address certifiable autonomy: selecting a set of model hyper-parameters $(\theta,\lambda)$ that achieves multiple user-specified levels of risk. As shown in Fig. 2, we formalize this problem by considering a joint distribution $\mathcal{D}$ over scenarios $\mathcal{\xi}:=(e,q,l)$, where $e$ is an environment (a POMDP) with dynamics $f_{e}$, $q$ is a stochastic function describing the human’s intent and partially observed through $f$, and $l$ is a cost function encoding the robot’s goal, which is assumed to be known a-priori by both the human and robot. We do not assume knowledge of $\mathcal{D}$ except for the availability of a modestly-sized calibration dataset $\mathcal{C}$ containing $500$ samples from $\mathcal{D}$. We formalize certifiable autonomy in our context as (i) risk calibration: the robot must meet a set of user-specified risk levels $(R_{1},\dotsc,R_{K})$ with user-specified probability over new scenarios $\xi\sim\mathcal{D}$, and (ii) flexible autonomy: the policy $\Pi^{R}$ should return a set of model hyper-parameters that control each risk but allow different high-level behaviors.
@@ IV-A Background: Statistical Risk Calibration
Assume that we are given a risk signal $R$ that we wish to control, where $R\in[0,1]$ measures an expected loss as a function of the prediction threshold $\lambda$ and model hyperparameters $\theta$. Here, we let $\phi:=(\lambda,\theta)\in\Phi$ be the pair of prediction parameters we wish to optimize, where $\Phi=\Lambda\times\Theta$. For fixed parameters $\phi$, the expected loss is itself a function of the context, prediction sets, and true labels over the unknown distribution $\mathcal{D}$, i.e.
$R(\phi)=\mathbb{E}^{(\bar{x},z)\sim\mathcal{D}}\Big{[}L(\bar{x},\mathcal{T}_{\phi}(\bar{{x}}),z)\Big{]},$ (11)
@@
$R_{\text{cov}}(\phi)=\mathbb{P}^{(\bar{x},z)\sim\mathcal{D}}\big{(}u^{R*}(z^{*})\notin\mathcal{T}_{\phi}(\bar{{x}})\big{)}.$ (12)
@@
Calibrating the Predicted Action Set. We will assume access to a calibration set $\mathcal{C}=\{(\bar{x}_{i},z_{i})\}_{i=1}^{M}$ of i.i.d. random variables drawn from $\mathcal{D}$, which we will use to estimate the risks. We seek to take the (uncalibrated) prediction model $g_{\theta}^{*}:\mathcal{X}^{t}\times\mathcal{Z}\rightarrow\mathcal{[}0,1]$ that produces softmax scores for each intent-conditioned action $u^{R*}(z)$. As described in Sec. III-B, we post-process the raw model outputs in $[0,1]$ to generate a prediction set $\mathcal{T}_{\lambda,\theta}(\bar{x})$ containing actions; this set is parameterized by a low-dimensional set of parameters $\lambda\in\Lambda$ and $\theta\in\Theta$, where $\Lambda$ is a finite set of prediction thresholds values one wishes to test, and $\Theta$ is a finite set of model hyperparameters, such as temperature. Then, we use the calibration set in order to choose the parameter pair $(\lambda,\theta)$ to control a user-specified risk, regardless of the quality of the predictor $g_{\theta}^{*}$.
Ahead of calibration, we set a desired risk threshold $\alpha$. Our goal is to identify a set $\hat{\Phi}\subseteq\Phi$ such that for any $\phi\in\hat{\Phi}$, $R(\phi)\leq\alpha$ with some user-defined probability $\delta$. In particular, the probability $\delta$ is with respect to the randomness in sampling over the calibration dataset $\mathcal{C}$, which itself is randomly sampled from the unknown distribution $\mathcal{D}$.
Multi-Hypothesis Testing for a Single Risk. Since the prediction set $\mathcal{T}_{\phi}$ is controlled by low-dimensional hyperparamters $\phi$ drawn from the set $\hat{\Phi}$, controlling a single risk is a multiple-hypothesis testing problem [5]. For each $j\in\{1,...,|\Phi|\}$, we consider the hypothesis $\mathcal{H}^{j}$ such that the risk $R(\phi^{j})$ is not controlled, where $\phi^{j}\in\Phi$. Therefore, rejecting $\mathcal{H}^{j}$ is equivalent to certifying that the risk is controlled. For a calibration set size $M$, define the empirical risk estimate on the calibration set:
$\hat{R}^{j}=M^{-1}\sum_{i=1}^{M}L(\bar{x}_{i},\mathcal{T}_{\phi^{j}}(\bar{x}_{i}),z_{i}).$ (13)
@@
We now have left to construct our set $\hat{\Phi}$ of low-dimensional parameters $\phi^{j}$ that reject $\mathcal{H}^{j}$ and control the risk $R$. Bounding (11) for all $\phi\in\hat{\Phi}$ requires that the $p$-values hold simultaneously; any nontrivial subset $\hat{\Phi}\subseteq\Phi$ that controls the risk is said to control the family-wise error rate (FWER). A simple but powerful approach, which we use in the following analysis, is to apply a union bound over a coarse grid $\mathcal{J}$ of initializations (e.g. each item in $\mathcal{J}$ is an equally spaced grid of indices of $\Phi$ ) in an iterative procedure called fixed-sequence testing [43, 5]. In fixed-sequence testing, for each $j\in\mathcal{J}$, the set $\hat{\Phi}$ of valid prediction thresholds is initialized as the empty set and grown according to the rule
$\hat{\Phi}\leftarrow\begin{cases}\hat{\Phi}\cup\{\phi^{l}\}\quad&\phi^{l}\notin\hat{\Phi}\text{ and }p^{l}\leq\delta/|\mathcal{J}|,\quad l\geq j\quad j\in[\mathcal{J}]\\ \hat{\Phi}&o.w.\end{cases}$ (15)
That is, parameters $\phi^{j}$ are only added to $\hat{\Phi}$ if $\mathcal{H}^{j}$ is rejected, eliminating the need for a union bound over a large set of parameters. The set of parameters that satisfy the risk bound is given by
$\hat{\Phi}:=\{\phi^{j}:p^{j}\leq\delta/|\mathcal{J}|\}.$ (16)
@@
$\mathbb{P}^{(\bar{x},z)\sim\mathcal{\mathcal{missing}}{D}^{M}}\Bigg{(}\sup_{\phi\in\hat{\Phi}}\big{\{}R(\phi)\big{\}}\leq\alpha\Bigg{)}\geq 1-\delta,$ (17)
where the supremum over the empty set is defined as $\infty$- ∞. The calibration procedure thus yields $\hat{\Phi}$, which is a set of values $\phi^{j}$ that each control the risk $R(\phi^{j})$ to the desired level $\alpha$ (with probability $1-\delta$ over the randomness in the calibration dataset).
@@ IV-B Single-Step, Single-Risk Control
Consider a single-step setting ($T=1$) where we use risk calibration parameters $\phi\in\hat{\Phi}$ to generate predicted action sets and seek help whenever the prediction set is not a singleton (cf. Sec. III-B). If the FWER-controlling parameter set $\hat{\Phi}$ is non-empty, then with probability $1-\delta$ over the sampling of the calibration set, the new scenarios drawn from $\mathcal{D}$ incur at most $\alpha_{1}$ rate of optimal action miscoverage.
@@
The proof follows immediately from application of fixed-sequence testing to the $p$-values obtained from the Hoeffding-Bentkus inequality, as given in [5], and is identical to the conformal prediction setting [36]. ∎
@@ IV-C Single-Step, Multi-Risk Control
$R_{\text{help}}(\phi)=\mathbb{P}^{(\bar{x},z)\sim\mathcal{D}}\big{(}|\mathcal{T}_{\phi}(\bar{{x}})|>1\big{)}.$ (18)
@@
$\mathcal{H}^{j}_{k}:\quad R_{k}(\phi^{j})\geq\alpha_{k}\quad k\in\{\textnormal{cov},\ \textnormal{help}\}$ (19)
@@
Consider a single-step setting where we use risk calibration parameters $\phi\in\hat{\Phi}$ to generate prediction sets and seek help whenever the prediction set is not a singleton. Let the upper bound on the help rate (18) be set to $\alpha_{2}$. If the FWER-controlling parameter set $\hat{\Phi}$ is non-empty, then with probability $1-\delta$ over the sampling of the calibration set, the new scenarios drawn from $\mathcal{D}$ incur at most $\alpha_{1}$ rate of optimal action miscoverage and at most $\alpha_{2}$ rate of human help.
@@ IV-D Multi-Step, Single-Risk Control
Sequence-Level Risk Calibration. Similar to [36], the key idea is to (i) lift the data to sequences and (ii) perform the LTT procedure using a carefully designed score function that allows for causal reconstruction of the prediction set at test time. We now consider a distribution $\bar{\mathcal{D}}$ of lifted contexts induced by $\mathcal{D}$, where each lifted context contains a state-intent pair $(\tilde{x},\bar{z})\sim{\bar{\mathcal{D}}}$. The lifted state-intent pairs are given as $\tilde{x}=(\bar{x}_{1},...,\bar{x}_{T})$ and $\bar{z}=(z_{1},...,z_{T})$ respectively. Here, $\tilde{x}$ arises from the robot having performed the correct action in previous steps. Using the robot policy specified in Section IIIB, there are three cases to consider: the robot will (i) take the only available (and optimal) action if $\mathcal{T}_{\phi}(\tilde{{x}})$ is a singleton, or (ii) ask for clarification of the human’s intent $z$ if the action set $\mathcal{T}_{\phi}(\tilde{{x}})$ is not a singleton. We bound the risk associated with case (iii): the optimal action is not in the prediction set as follows. Let $\bar{u}^{R*}(\bar{z}):=(u^{R*}(z_{1}),...,u^{R*}(z_{T}))$ denote the sequence of optimal robot actions. Let the sequence-level confidence be given as the lowest confidence over the timesteps
@@
where the corresponding sequence-level prediction set is given as $\bar{\mathcal{T}}_{\phi}(\tilde{x})=\{\bar{u}\in\mathcal{U}^{T}:\exists\bar{z}\in\mathcal{Z}^{T}\;\text{ s.t. }\bar{u}=\bar{u}^{R*}(\bar{z})\text{ and }\bar{g}_{\theta}^{*}(\tilde{x},\bar{z})\geq\lambda)\}$.
Causal Reconstruction of Sequence-Level $\bar{\mathcal{T}}_{\phi}$. The sequence-level prediction set $\bar{\mathcal{T}}_{\phi}$ is constructed with the full sequence $\bar{z}$ as labels, which depend causally on the sequence $\bar{x}$. Hence, we do not have the entire sequence $\bar{z}$ a-priori; the robot must instead construct the prediction set at each time-step in a causal manner (i.e., relying only on current and past observations). Let $\mathcal{T}_{\phi}^{t}(\bar{x}_{t}):=\{u\in\mathcal{U}:\exists z\in\mathcal{Z}\;\text{ s.t. }u=u^{R*}(z)\text{ and }g_{\theta}^{*}(\bar{x}_{t},z_{t})\geq\lambda)\}$ be the instantaneous action prediction set at time $t$. We construct $\bar{\mathcal{T}}_{\phi}$ in a causal manner using
${\mathcal{T}}_{\phi}(\tilde{x}):=\mathcal{T}_{\phi}^{1}(\bar{x}_{1})\times\mathcal{T}_{\phi}^{2}(\bar{x}_{2})\times...\times\mathcal{T}_{\phi}^{T-1}(\bar{x}_{T-1}).$ (21)
@@
Consider a multi-step setting where we use risk calibration parameters $\phi\in\hat{\Phi}$ and the sequence-level confidence (20) to generate sequence-level prediction sets and seek help whenever the prediction set is not a singleton. If the FWER-controlling parameter set $\hat{\Phi}$ is non-empty, then with probability $1-\delta$ over the sampling of the calibration set, the new scenarios drawn from $\bar{\mathcal{D}}$ under $\Pi^{R}$ and using the causally reconstructed predicted action set (21) incur at most $\alpha_{1}$ rate of action miscoverage.
@@
Let $\phi\in\hat{\Phi}$, where $\hat{\Phi}$ controls the sequence-level FWER for the non-causal set $\bar{\mathcal{T}}_{\phi}(\tilde{x})$ at level $\alpha_{1}$. We first show that $\bar{u}^{R*}(\bar{z})\in\bar{\mathcal{T}}_{\phi}(\tilde{x})\Longleftrightarrow\bar{u}^{R*}(\bar{z})\in{\mathcal{T}}_{\phi}(\tilde{x})$. For any $\bar{z}\in\bar{\mathcal{T}}_{\phi}(\tilde{x})$,
$\bar{u}^{R*}(\bar{z})\in\bar{\mathcal{T}}_{\phi}(\tilde{x})$ $\Longleftrightarrow\min_{t\in[T]}\;\;g_{\theta}^{*}(\bar{x}_{t},z_{t})\geq\lambda$ (22)
@@
$\Longleftrightarrow u^{R*}(z_{t})\in{\mathcal{T}}_{\phi}^{t}(\bar{x}_{t})\quad\forall t$
$\Longleftrightarrow\bar{u}^{R*}(\bar{z})\in{\mathcal{T}}_{\phi}(\tilde{x}).$

### v2

@@ III-C Goal: Certifiable Autonomy
Situational ambiguity results in many potentially correct robot actions arising with potentially no safe external resolution, save for direct human intervention (see e.g. Fig. 1). Our goal in this work is to address certifiable autonomy: selecting a set of model parameters $(\theta,\lambda)$ that achieve multiple user-specified levels of risk. As shown in Fig. 2, we formalize this problem by considering a joint distribution $\mathcal{D}$ over scenarios $\mathcal{\xi}:=(e,H,l,h)$, where $e$ is an environment (a POMDP) with dynamics $f$ are parameterized by $e$, $H$ is a particular human with stochastic intent dynamics $q$, and $l$ and $h$ are respectively a cost and constraint function encoding the robot’s goal, which is assumed to be known a-priori by both the human and robot. We do not assume knowledge of $\mathcal{D}$ except for the availability of a modestly-sized calibration dataset $D_{\text{cal}}$ containing $400$ samples from $\mathcal{D}$. We formalize certifiable autonomy in our context as (i) risk calibration: the robot must meet a set of user-specified risk levels $(\alpha_{1},\dotsc,\alpha_{K})$ with user-specified probability over new scenarios $\xi\sim\mathcal{D}$, and (ii) flexible autonomy: the policy $\Pi^{R}$ should return a set of model parameters that control each risk but allow different high-level behaviors.
@@ IV-A Background: Statistical Risk Calibration
Assume that we are given a risk signal $R$ that we wish to control, where $R\in[0,1]$ measures an expected loss as a function of the prediction threshold $\lambda$ and model parameters $\theta$. For fixed parameters $(\lambda,\theta)$, the expected loss is itself a function of the context, prediction sets, and true labels over the unknown distribution $\mathcal{D}$, i.e.
$R(\lambda,\theta)=\mathbb{E}^{(\bar{x},z)\sim\mathcal{D}}\Big{[}L(\bar{x},\mathcal{T}_{\lambda,\theta}(\bar{{x}}),z)\Big{]},$ (12)
@@
$R_{\text{cov}}(\lambda,\theta)=\mathbb{P}^{(\bar{x},z)\sim\mathcal{D}}\big{(}u^{R*}(z^{*})\notin\mathcal{T}_{\lambda,\theta}(\bar{{x}})\big{)}.$ (13)
@@
Calibrating the Predicted Action Set. We will assume access to a calibration set $D_{\text{cal}}=\{(\bar{x}_{i},z_{i})\}_{i=1}^{M}$ of i.i.d. random variables drawn from $\mathcal{D}$, which we will use to estimate the risks. We seek to take the (uncalibrated) prediction model $g_{\theta}^{*}:\mathcal{X}^{t}\times\mathcal{Z}\rightarrow\mathcal{[}0,1]$ that produces softmax scores for each intent-conditioned action $u^{R*}(z)$. As described in Sec. III-B, we post-process the raw model outputs in $[0,1]$ to generate a prediction set $\mathcal{T}_{\lambda,\theta}(\bar{x})$ containing actions; this set is parameterized by a low-dimensional set of parameters $\lambda\in\Lambda$ and $\theta\in\Theta$, where $\Lambda$ is a finite set of prediction thresholds values one wishes to test, and $\Theta$ is a finite set of model parameters, such as temperature. Then, we use the calibration set in order to choose the parameter pair $(\lambda,\theta)\in\Lambda\times\Theta$ to control a user-specified risk, regardless of the quality of the predictor $g_{\theta}^{*}$. In future sections, we let $\Phi$ denote this product search space.
Ahead of calibration, we set a desired risk threshold $\alpha$. Our goal is to identify a set $\hat{\Phi}\subseteq\Phi$ such that for any $(\lambda,\theta)\in\hat{\Phi}$, $R(\lambda,\theta)>\alpha$ (i.e., the risk is not controlled) with some user-defined probability $\delta$. In particular, the probability $\delta$ is with respect to the randomness in sampling over the calibration dataset $D_{\text{cal}}$, which itself is randomly sampled from the unknown distribution $\mathcal{D}$.
Multi-Hypothesis Testing for a Single Risk. Since the prediction set $\mathcal{T}_{\lambda,\theta}$ is controlled by low-dimensional hyperparamters $(\lambda,\theta)$ drawn from the set $\hat{\Phi}$, controlling a single risk is a multiple-hypothesis testing problem [5]. For each $j\in\{1,...,|\Phi|\}$, we consider the hypothesis $\mathcal{H}^{j}$ such that the risk $R(\lambda^{j},\theta^{j})$ is not controlled, where $\lambda^{j},\theta^{j}\in\Phi$. Therefore, rejecting $\mathcal{H}^{j}$ is equivalent to certifying that the risk is controlled. For a calibration set size $M$, define the empirical risk estimate on the calibration set:
$\hat{R}^{j}=\frac{1}{M}\sum_{i=1}^{M}\;L(\bar{x}_{i},\mathcal{T}_{\lambda^{j},\theta^{j}}(\bar{x}_{i}),z_{i}).$ (14)
@@
We now have left to construct our set $\hat{\Phi}$ of low-dimensional parameters $(\lambda^{j},\theta^{j})$ that reject $\mathcal{H}^{j}$ and control the risk $R$. Bounding (12) for all $(\lambda,\theta)\in\hat{\Phi}$ requires that the $p$-values hold simultaneously; any nontrivial subset $\hat{\Phi}\subseteq\Phi$ that controls the risk is said to control the family-wise error rate (FWER). A simple but powerful approach, which we use in the following analysis, is to apply a union bound over a coarse grid $\mathcal{J}$ of initializations (e.g. each item in $\mathcal{J}$ is an equally spaced grid of indices of $\Phi$ ) in an iterative procedure called fixed-sequence testing [46, 5]. In fixed-sequence testing, for each $j\in\mathcal{J}$, the set $\hat{\Phi}$ of valid prediction thresholds is initialized as the empty set and grown according to the rule
$\hat{\Phi}\leftarrow\begin{cases}\hat{\Phi}\cup\{(\lambda^{l},\theta^{l})\}\quad&(\lambda^{l},\theta^{l})\notin\hat{\Phi}\text{ and }p^{l}\leq\delta/|\mathcal{J}|,\\ \hat{\Phi}&o.w.\end{cases}$ (16)
for $l\geq j$ and $j\in[\mathcal{J}]$. That is, parameters $(\lambda^{j},\theta^{j})$ are only added to $\hat{\Phi}$ if $\mathcal{H}^{j}$ is rejected, eliminating the need for a union bound over a large set of parameters. The set of parameters that satisfy the risk bound is given by
$\hat{\Phi}:=\{\lambda^{j},\theta^{j}:p^{j}\leq\delta/|\mathcal{J}|\}.$ (17)
@@
$\mathbb{P}^{(\bar{x},z)\sim\mathcal{\mathcal{missing}}{D}^{M}}\Bigg{(}\sup_{(\lambda,\theta)\in\hat{\Phi}}\big{\{}R(\lambda,\theta)\big{\}}\leq\alpha\Bigg{)}\geq 1-\delta,$ (18)
where the supremum over the empty set is defined as $\infty$- ∞. The calibration procedure thus yields $\hat{\Phi}$, which is a set of values $\lambda^{j},\theta^{j}$ that each control the risk $R(\lambda^{j},\theta^{j})$ to the desired level $\alpha$ (with probability $1-\delta$ over the randomness in the calibration dataset).
@@ IV-B Single-Step, Single-Risk Control
Consider a single-step setting ($T=1$) where we use risk calibration parameters $(\lambda,\theta)\in\hat{\Phi}$ to generate predicted action sets and seek help whenever the prediction set is not a singleton (cf. Sec. III-B). If the FWER-controlling parameter set $\hat{\Phi}$ is non-empty, then with probability $1-\delta$ over the sampling of the calibration set, the new scenarios drawn from $\mathcal{D}$ incur at most $\alpha_{1}$ rate of optimal action miscoverage.
@@
The proof follows immediately from application of fixed-sequence testing to the $p$-values obtained from the Hoeffding-Bentkus inequality, as given in [5], and is nearly identical to the conformal prediction setting [36] up to the choice of concentration inequality. ∎
@@ IV-C Single-Step, Multi-Risk Control
$R_{\text{help}}(\lambda,\theta)=\mathbb{P}^{(\bar{x},z)\sim\mathcal{D}}\big{(}|\mathcal{T}_{\lambda,\theta}(\bar{{x}})|>1\big{)}.$ (19)
@@
$\mathcal{H}^{j}_{k}:\quad R_{k}(\lambda^{j},\theta^{j})\geq\alpha_{k}\quad k\in\{\textnormal{cov},\ \textnormal{help}\}$ (20)
@@
Consider a single-step setting where we use risk calibration parameters $(\lambda,\theta)\in\hat{\Phi}$ to generate prediction sets and seek help whenever the prediction set is not a singleton. Let the upper bound on the help rate (19) be set to $\alpha_{2}$. If the FWER-controlling parameter set $\hat{\Phi}$ is non-empty, then with probability $1-\delta$ over the sampling of the calibration set, the new scenarios drawn from $\mathcal{D}$ incur at most $\alpha_{1}$ rate of optimal action miscoverage and at most $\alpha_{2}$ rate of human help.
@@ IV-D Multi-Step, Single-Risk Control
Sequence-Level Risk Calibration. Similar to [36], the key idea is to (i) lift the data to sequences and (ii) perform the LTT procedure using a carefully designed score function that allows for causal reconstruction of the prediction set at test time. We now consider a distribution $\bar{\mathcal{D}}$ of lifted contexts induced by $\mathcal{D}$, where each lifted context contains a state-intent pair $(\tilde{x},\bar{z})\sim{\bar{\mathcal{D}}}$. The lifted state-intent pairs are given as $\tilde{x}=(\bar{x}_{1},...,\bar{x}_{T})$ and $\bar{z}=(z_{1},...,z_{T})$ respectively. Here, $\tilde{x}$ arises from the robot having performed the correct action in previous steps. Using the robot policy specified in Section IIIB, there are three cases to consider: the robot will (i) take the only available (and optimal) action if $\mathcal{T}_{\lambda,\theta}(\tilde{{x}})$ is a singleton, or (ii) ask for clarification of the human’s intent $z$ if the action set $\mathcal{T}_{\lambda,\theta}(\tilde{{x}})$ is not a singleton. We bound the risk associated with case (iii): the optimal action is not in the prediction set as follows. Let $\bar{u}^{R*}(\bar{z}):=(u^{R*}(z_{1}),...,u^{R*}(z_{T}))$ denote the sequence of optimal robot actions. Let the sequence-level confidence be given as the lowest confidence over the timesteps
@@
where the corresponding sequence-level prediction set is given as $\bar{\mathcal{T}}_{\lambda,\theta}(\tilde{x})=\{\bar{u}\in\mathcal{U}^{T}:\exists\bar{z}\in\mathcal{Z}^{T}\;\text{ s.t. }\bar{u}=\bar{u}^{R*}(\bar{z})\text{ and }\bar{g}_{\theta}^{*}(\tilde{x},\bar{z})\geq\lambda)\}$.
Causal Reconstruction of Sequence-Level $\bar{\mathcal{T}}_{\lambda,\theta}$. The sequence-level prediction set $\bar{\mathcal{T}}_{\lambda,\theta}$ is constructed with the full sequence $\bar{z}$ as labels, which depend causally on the sequence $\bar{x}$. Hence, we do not have the entire sequence $\bar{z}$ a-priori; the robot must instead construct the prediction set at each time-step in a causal manner (i.e., relying only on current and past observations). Let $\mathcal{T}_{\lambda,\theta}^{t}(\bar{x}_{t}):=\{u\in\mathcal{U}:\exists z\in\mathcal{Z}\;\text{ s.t. }u=u^{R*}(z)\text{ and }g_{\theta}^{*}(\bar{x}_{t},z_{t})\geq\lambda)\}$ be the instantaneous action prediction set at time $t$. We construct $\bar{\mathcal{T}}_{\lambda,\theta}$ in a causal manner using
${\mathcal{T}}_{\lambda,\theta}(\tilde{x}):=\mathcal{T}_{\lambda,\theta}^{1}(\bar{x}_{1})\times\mathcal{T}_{\lambda,\theta}^{2}(\bar{x}_{2})\times...\times\mathcal{T}_{\lambda,\theta}^{T-1}(\bar{x}_{T-1}).$ (22)
@@
Consider a multi-step setting where we use risk calibration parameters $(\lambda,\theta)\in\hat{\Phi}$ and the sequence-level confidence (21) to generate sequence-level prediction sets and seek help whenever the prediction set is not a singleton. If the FWER-controlling parameter set $\hat{\Phi}$ is non-empty, then with probability $1-\delta$ over the sampling of the calibration set, the new scenarios drawn from $\bar{\mathcal{D}}$ under $\Pi^{R}$ and using the causally reconstructed predicted action set (22) incur at most $\alpha_{1}$ rate of action miscoverage.
@@
Let $(\lambda,\theta)\in\hat{\Phi}$, where $\hat{\Phi}$ controls the sequence-level FWER for the non-causal set $\bar{\mathcal{T}}_{\lambda,\theta}(\tilde{x})$ at level $\alpha_{1}$. We first show that $\bar{u}^{R*}(\bar{z})\in\bar{\mathcal{T}}_{\lambda,\theta}(\tilde{x})\Longleftrightarrow\bar{u}^{R*}(\bar{z})\in{\mathcal{T}}_{\lambda,\theta}(\tilde{x})$. For any $\bar{z}\in\bar{\mathcal{T}}_{\lambda,\theta}(\tilde{x})$,
$\bar{u}^{R*}(\bar{z})\in\bar{\mathcal{T}}_{\lambda,\theta}(\tilde{x})$ $\Longleftrightarrow\min_{t\in[T]}\;\;g_{\theta}^{*}(\bar{x}_{t},z_{t})\geq\lambda$ (23)
@@
$\Longleftrightarrow u^{R*}(z_{t})\in{\mathcal{T}}_{\lambda,\theta}^{t}(\bar{x}_{t})\quad\forall t$
$\Longleftrightarrow\bar{u}^{R*}(\bar{z})\in{\mathcal{T}}_{\lambda,\theta}(\tilde{x}).$